OpenStudy (anonymous):

Find the linear approximation of f(x)=√4-x at a=0 and use it to approximate √3.9 and √4.1. L(x)= √3.9≈ √4.1≈

myininaya (myininaya):

So we just need to find the tangent line to f at a=0 We need the following info f'(x) f'(a) Then enter into here: y-f(a)=f'(a)(x-a)

myininaya (myininaya):

We will use that y to approximate f=sqrt(4-x)

OpenStudy (anonymous):

so f'(x) is (1/2) (4-x)^(-1/2) * (-1) right?

OpenStudy (anonymous):

f'(a) = (1/2)(4-a)^(-1/2) * (-1) right?

myininaya (myininaya):

yep...and a=0, correct? So f'(0)=-1/2 * (4-0)^(-1/2)